OpenStudy (anonymous):

Which property justifies this statement? "If 7 = x + 4, then x + 4 = 7." A. commutative B. symmetric C. associative D. transitive

OpenStudy (anonymous):

symmetric if left = right then right =left i would say

OpenStudy (anonymous):

?

OpenStudy (freckles):

commutative is talking about the order in an expression that is this property looks like this: a+b = b+a (the order in the expression changes) symmetric is talking about all one direction equality implies the other direction of equality that is this property looks like this: a=b implies b=a associative is talking about the grouping changing in the expression that is this property looks like this: (a+b)+c = a+(b+c) transitive is talking about how two different certain equalities imply another that is this property looks like this: a=b and b=c implies a=c

OpenStudy (anonymous):

SO symmetric

OpenStudy (freckles):

yeah

OpenStudy (anonymous):

I would go with symmetric. :)

OpenStudy (freckles):

that is what @jamaica25 was trying to say she was calling the left hand expression of the equality left and the right hand expression of the equality right that is we have 7=x+4 so 7 is the left hand expresssion (she called just left) and x+4 is the right hand expression (she called right) so she said left=right implies right=left which is the exactly the symmetric property

OpenStudy (anonymous):

O